OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle, whose legs have lengths of 12 meters and 35 meters?

OpenStudy (anonymous):

Use the pythagorean theorem  In your case, b = 12 m and c = 35 m, the hypotenuse a is: a^2 = b^2 + c^2 a^2 = 12^2 + 35^2 a = sqrt(12^2 + 35^2) a = ...?